OpenStudy (goformit100):

The least number that is divisible by all the numbers from 1 to 10 (both inclusive) is (A) 10 (B) 100 (C) 504 (D) 2520

OpenStudy (goformit100):

@dan815

OpenStudy (dan815):

2,9 ,3,5 ar prime in here so

OpenStudy (goformit100):

then ?

OpenStudy (dan815):

no sry thats wrong lemme think xD

OpenStudy (dan815):

it is 2520

OpenStudy (anonymous):

The required number is LCM of 1,2,3,4,5,6,7,8,9,10 LCM= 2*2*2*5*3*3*7 =2520

OpenStudy (goformit100):

ok then ?

OpenStudy (dan815):

as 2+5+2+0 = 9 so 3 goes in it and 5 goes it in it

OpenStudy (goformit100):

Thanks

OpenStudy (dan815):

now u need to see if 7 also goes in it and ur proof is complete

OpenStudy (dan815):

forgot the rule for 7 divisibility

OpenStudy (dan815):

@Claira16 hey how did u get that LCM formula

OpenStudy (dan815):

did u look at numbers 8 5 9 and 7 10 2*2*2*5*3*3*7

OpenStudy (dan815):

ok i see cool